a 700 ml sample of gas has a pressure of 760.0 mm Hg at a temperature of 21.5 C.if the temperature of the gas increases to 30.0 C and the volume of the gas is compressed to 200.0 mL, what is the new pressure in kPa


1
Expert's answer
2021-02-24T03:42:00-0500

There are two stages of this gas


1st Stage:

"V_1" =700 ml

"P_1" =760mm Hg

"T_1" =21.5 degre celsius = 294.5 kelvin


Second stage:

"V_2" =200 ml

"T_2" =30 degre cecsius = 303 kelvin

"P_2=P_2" (let)

From the gas eqn ,

PV=nRT

n="\dfrac{PV}{RT}"


Now we know that in both the stages moles are conserved

"\Rightarrow" "n_1=n_2"

"\Rightarrow" "\dfrac{P_1V_1}{RT_1}" = "\dfrac{P_2V_2}{RT_2}"

as R is constant from both sides, So finally


"P_2" ="\dfrac{P_1V_1T_2}{T_1V_2}"


"P_2" = "\dfrac{760\times700\times303}{294.5\times200}" = 2736.77 mm Hg= 364.87 KPa